OpenStudy (anonymous):

3x+5(y+2)=1 8y=-3x Solve

OpenStudy (anonymous):

Simultaneous equations. \[3x+5y=-9 (1)\] \[8y+3x=0(2)\] \[(2)-(1)\] the 3x cancels out. "--9" becomes +9. Negative Negative equals Positive. \[3y= 9\]

OpenStudy (anonymous):

Now solve for y. Once you solve y. You can then solve x by subbing the y component into one of the two equations.

OpenStudy (anonymous):

The equation reads 3x+5(y+2)=1 How did u get3x+5=-9(1) I am so confused

OpenStudy (anonymous):

Please explain. I am thankful for answer but i want to understand how you got 3x+5y=-9(1)

OpenStudy (anonymous):

Expand "5(y+2)"

OpenStudy (anonymous):

5y+10

OpenStudy (anonymous):

That will give you 5y+10 Then the whole thing will be 3x+ 5y + 10. Good.

OpenStudy (anonymous):

So now you move the +10 to the Right Hand Side. You will then have to change the sign of the "+10" to "-10". The equation will be 3x+5y= 1 - 10

OpenStudy (anonymous):

Ok so u minus -10 from both sides so you get -9

OpenStudy (anonymous):

Yes. Well Done.

OpenStudy (anonymous):

My teacher is confusing me bc i understand that part but then after i do the first part he wants me to cancel out both 5y and the 8y

OpenStudy (anonymous):

How do the 3x cancel out if they are both positive

OpenStudy (anonymous):

I Am sorry but please explain to me. I understand but i dont see why you have a 1 next to the (9) and a 2 after (0)

OpenStudy (anonymous):

Subtract the first equation by the second equation.

OpenStudy (anonymous):

So the LHS(Left Hand Side) of the first equation is subtracted by the LHS of the second equation. Same goes for the RHS(Right Hand Side)

OpenStudy (anonymous):

3x+5y=-9 3x8y=0

OpenStudy (anonymous):

SO \[(8y+3x)-(3x+5y) = 0-(-9)\]

OpenStudy (anonymous):

So I'm Subtracting the FIRST equation BY the SECOND Equation.

OpenStudy (anonymous):

YOu can SUbtract the SECOND EQUATION with the FIRST EQUATION but that will leave you with minus y's all over the place. It's best to get positive numbers than negative numbers but you get the same answer with boths ways.

OpenStudy (anonymous):

both ways*

OpenStudy (anonymous):

Omg thank you so much. I couldnt figure it out. I got it

OpenStudy (anonymous):

Okay. That's good to hear. If you need further help, you're welcome to ask about it again. Hope everything is clear with that. Good luck!

OpenStudy (anonymous):

Thanks